OpenStudy (anonymous):

Tom and Jerry are jogging in the woods along straight, level paths that intersect at right angles. Tom is 3 miles from the intersection and jogging away from it at the rate of 2 miles per hour. Jerry is 4 miles from the intersection and is jogging away from it at the rate of 3 miles per hour. At what rate is the distance between Tom and Jerry increasing?

OpenStudy (anonymous):

3.6 miles per hour. Care for an explanation? hehe!

OpenStudy (anonymous):

yeah please

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

\(z^2 = x^2 + y^2\) \( 2z\dfrac{dz}{dt} = 2x\dfrac{dx}{dt} + 2y\dfrac{dy}{dt} \) \( z\dfrac{dz}{dt} = x\dfrac{dx}{dt} + y\dfrac{dy}{dt} \) \( \dfrac{dz}{dt} = \dfrac{ x \dfrac{dx}{dt} + y \dfrac{dy}{dt}}{z} ~~~~~(1)\) Plug in the values below into equation (1) above to find the answer. \( x = 3, ~\dfrac{dx}{dt} = 2, ~y = 4, ~\dfrac{dy}{dt} = 3\) \(z = \sqrt{3^2 + 4^2} = 5\)

OpenStudy (anonymous):

Thank you that was a awesome respond

OpenStudy (mathstudent55):

You're welcome.